OpenStudy (anonymous):

find the center of curvature for the point P (2,0) for y=ln (x-1)

OpenStudy (anonymous):

What does the center of curvature mean?

OpenStudy (amistre64):

it means, i believe find the center of the circle that is tangent to (2,0) that has the same curvature

OpenStudy (amistre64):

http://mathworld.wolfram.com/RadiusofCurvature.html k is the usual name for curvature; the radius is 1/k k = y''/(1+(y')^2)^(3/2) y=ln(x-1) y' = (x-1)^(-1) y'' = -(x-1)^(-2) \[k=-\frac{1}{(x-1)^2}*\frac{1}{\sqrt[2]{(1+(x-1)^2)}^3}\] \[k=-\frac{1}{(2-1)^2}*\frac{1}{\sqrt[2]{(1+(2-1)^2)}^3}\] \[k=-\frac{1}{\sqrt[2]{2}^3}\] \[k=-\frac{1}{\sqrt[2]{8}}\] \[k=-\frac{1}{2\sqrt{2}}\] \[R=\frac{1}{k}=-2\sqrt{2}\]

OpenStudy (amistre64):

R is spose to be an absolute value so ignore the negative :) now if we now the tangent to this curve at the point; then we know that "normal" (th eperp to the tangent) at this point as well y' = (x-1)^(-1); at x=2 we get a slope of 1 our normal to the curve is the slope -1. which is a 45 degree angle; the distance to the center is 2sqrt(2)

OpenStudy (amistre64):

im gonna say the center is at 2 down and 2 more over from (2,0) -> (4,2)